# Multi-Modal Framing Analysis of News

### v2

@@ 1 Introduction
Figure 1: News can be intentionally framed to affect reader perception. Editorial choices decide what is communicated through the words and the images. Our approach systematically detects this framing.
@@
While visual framing analysis has been explored in communication science and journalism studies (Wessler et al., 2016; Powell et al., 2017), existing studies doing computational framing analysis have largely ignored this crucial aspect (Ali and Hassan, 2022). Further, they have primarily focused on a fixed set of labels for framing analysis, performing prediction in a multi-class setting. This substantially limits the information one can derive from predictions as an article can convey several frames (Figure 1) and fine-grained analysis of framing within a topic necessitates frames specific to a particular issue. Large vision and language models are particularly well suited for conducting this task at scale, considering the semantic understanding embedded into them through large scale pre-training. Our contributions include the following.
@@
  • We provide a large-scale dataset  [^1] https://huggingface.co/datasets/copenlu/mm-framing of 500k U.S.-based news articles for framing analysis, automatically labeled with generic and issue-specific frames, validated through a smaller set of human annotations.
@@ 2 News Framing
2.1 Framing through visuals
@@
Photographs are an important vehicle of framing as a reader may process textual and visual messages differently. Readers may focus on photographs without also reading an accompanying story Miller (1975) or might select which news story to read depending on the image thumbnail. Images are potent framing tools because they evoke immediate emotional responses, provide contextual cues, and sometimes contradict textual narratives Geise and Xu (2024). For visual data, however, automation is more challenging because computational image analysis often struggles with connotative and symbolic elements that are readily discernible to human annotators Rodriguez and Dimitrova (2011). While some recent advancements in machine learning and computer vision (e.g., object detection, facial recognition, sentiment analysis on images) offer promising avenues for automated visual framing analysis, the complexity of symbolic and ideological meanings typically requires human interpretation Coleman (2010).
Understanding multimodal framing is important for several reasons: it allows for a more granular examination of media bias, as textual analysis alone may miss the ideological and emotional undertones of visual elements Wessler et al. (2016); Geise et al. (2025). It improves fact-checking and detection of misleading content by identifying inconsistencies between textual claims and their supporting images. It also finds applications in political communication, journalism studies, and public policy analysis through tracing changes in framing over time and across outlets Baumgartner et al. (2008).
2.2 Integrative Framing Analysis
Integrative framing analysis is when both images and text are observed separately, but the results are integrated to form a more complete picture of framing analysis (Dan, 2017). Although there is broad agreement that visual and verbal elements should be studied side by side (Coleman, 2010; Rose, 2022), relatively few studies have effectively combined the two. Researchers often focus on textual framing Matthes and Kohring (2008) or visual framing Rodriguez and Dimitrova (2011) as separate domains, even though media messages typically blend both. Geise et al. (2013) and Coleman (2010) emphasize that analyzing visuals in isolation can obscure how words shape viewers’ interpretations—and vice versa. As computational tools become increasingly sophisticated and as scholarly interest in visual framing continues to grow, integrative approaches are poised to become more common. By integrating these modalities, scholars can better observe whether text and visuals reinforce each other or present divergent frames.
2.3 Automated Framing Analysis
@@
In computational studies, framing has been studied with the help of machine learning methods for content analysis at scale.  Ali and Hassan (2022); Vallejo et al. (2023); Otmakhova et al. (2024) survey these efforts towards on computational framing analysis, focusing on different aspects like methods, varying conceptualisations and inter-disciplinary connections of framings. The most widely used resource is one by  Card et al. (2015), who present the Media Frames Corpus (MFC), a dataset of US based news headlines annotated for generic frames. Most studies use it in a supervised, multi-class prediction setting with MFC frames for analysis of social networks (Mendelsohn et al., 2021) or discussion forums (Hartmann et al., 2019). There are also English-only and multi-lingual SemEval tasks on frame detection. (Piskorski et al., 2023a, b; Sajwani et al., 2024). Unsupervised approaches to framing analysis use framing lexica (Field et al., 2018), clustering (Burscher et al., 2016; Ajjour et al., 2019), and topic models (Nguyen, 2015). All of these approaches have focused on texts alone.
@@ 3 Dataset
We crawl news articles along with corresponding images from 28 US-based news agencies. We extract data for a 12 month period between May 2023 and April 2024 using the newsplease library (Hamborg et al., 2017). Our selected news sources reflect the entire political spectrum, based on data from AllSides Media Bias, [^2] https://www.allsides.com/media-bias/ratings an organization that assigns a rating of Left, Lean Left, Center, Lean Right, or Right to each media outlet. Our list of sources along with corresponding political leanings are listed in Appendix A. We query the publicly available Common Crawl archives [^3] https://commoncrawl.org/news-crawl for the corresponding publishers and extracted each article’s text, headline, publication date, image_urls and other metadata in JSON format. Post scraping, we filter extremely short and long articles, images with logos and other noise. Further details are provided in Appendix B. Our final dataset includes  500K articles and corresponding images. We show the distribution of the data across the time period of data collection in Figure 3.
@@ 4.1 Model Annotation
We use Large Language Models (LLM) and Vision-Language Models (VLM) to label several aspects of the text and image from news articles, respectively. In Table 1, we list the different aspects extracted per modality from the news articles. We also generate explanations for the decisions along each prediction for reducing hallucination and facilitating downstream qualitative analysis. For extracting generic frame across both modalities, we use the framing dimensions outlined by Boydstun et al. (2014), as provided in the Media Frames Corpus (Card et al., 2015), which includes 15 generic frames appropriate for analysis of US news. These frames are Economic, Capacity & Resources, Morality. Fairness & Equality, Legality, Constitutionality & Jurisprudence, Policy Prescription & Evaluation, Crime & Punishment, Security & Defense, Health & Safety, Quality of Life, Cultural Identity, Public Opinion, Political, External Regulation & Reputation and Other. The descriptions of each of the frames are provided in Appendix C. Differing from most previous studies on automated framing analysis, which assign a single frame to an article, we conduct the frame extraction in a multi-label setting, i.e., an article or an image can have more than one frame. This setting is much closer to the setup scholars use when conducting qualitative studies on framing (Dan, 2017). For the main subject of the articles, the models were instructed to extract entities central to the text or the image, if there is one.
Modality Aspect Description
Text Topic Broad topic of the article
Main subject Subject of the text
Generic Frame One or more of the 15 generic media frames
Issue Frame Inductive, open-ended framing of a topic
Image Caption Caption of the image
Main Subject Subject in the image
Generic Frame One or more of the 15 generic media frames
Table 1: List of aspects extracted per modality.
@@ 4.2 Models and Prompts
For text annotations, we use Mistral-7B. For images, we use Pixtral-12B (multi-modal model). We use the vLLM library for optimized memory usage and high-throughput inference, allowing us to conduct analysis on the entire dataset in a few days, demonstrating our approach’s scalability. See Appendix D for full modelling details.
We carefully craft prompts for the extraction of each aspect from the article. For extraction of frames expressed in the articles and images, we experiment with including definitions of framing and descriptions of each frame at several granularities. We benchmark these different prompts on a validation set. We also iteratively improve the prompt by looking at the category of errors and including those in the prompt. For instance, the Pixtral model has a tendency to predict the economic frame every time an entity in professional attire appears in the image, associating professional attire with being wealthy. We instruct the model to avoid such reliance. Finally, we report performance on a held-out test set. For issue-specific frames, we prompt the model by providing a definition of framing and some examples of issue-specific frames across topics. The task was then to analyse the article and generate an issue-specific frame with respect to the topic of the article. The full list of prompts for the extraction of each aspect is provided in Appendix  1 and  2.
@@ 4.3 Model Evaluation
To assess the quality of output frames from our approach, and compare against prior work, we evaluate our text frame extraction model on the existing large scale benchmark Media Frames Corpus (Card et al., 2015). This corpus provides frame labels by each annotator for over 32,000 articles from US news on topics like immigration, smoking, and same-sex marriage. We take the union of all frame labels assigned by the annotators and assign the up to the top 3 most frequent frames assigned by the annotators as the label for an article. We run our text framing analysis model on this dataset, providing the article of the text while allowing the model to generate multiple frames per article. To assess model performance, we calculate the intersection between the sets of model-annotated frames and human-annotated frames for each article in the dataset. 95.7% of the articles had a non-zero intersection, with at least one overlapping frame label, demonstrating that the model outputs accurate frames in most cases. Our model received a micro averaged F1 score of 0.5, with an averaged precision of 0.42 and averaged recall of 0.62. We include the scores per frame label in Appendix F. For prediction across 15 labels and a task as subjective as framing, we believe the model performs quite well. We further did qualitative assessment of model outputs by going through the model explanations for the predicted frames and found the model to be competent.
@@
For images, there is no existing benchmark with frame labels. Two of the authors of this study manually annotated 600 images for generic frames across a stratified (time-period and topics) sample of the dataset. We modify generic frame annotation guidelines released by Mendelsohn et al. (2021), adapting them to the image annotation setting. We set up a multi-label classification platform (as seen in Figure 12) where annotators are allowed to select one or more frames from the list given an image. Framing analysis is a highly subjective task, as is well established by prior work in communication studies as well as NLP. For images, the subjectivity increases substantially, given the limited amount of context available and requiring cultural and conceptual familiarity more so than needed for textual framing analysis (Geise and Baden, 2015). We then calculate the proportional overlap in frames between annotators, counting a non-zero intersection of frame annotations as 1 and zero intersections as 0. The proportional overlap between the annotators is 85%, demonstrating the high percentage of instances with overlapping frame labels. Using this dataset, we create an evaluation set for our vision model and get predictions from our pixtral model. Since ’Sports’ news are mostly labelled as ’cultural identity’, we remove them (8%of them) for our calculation. We take the union of the frame predictions for the two annotators and calculate the intersection with the model predictions for a particular image. The proportion of non-zero intersection instances between the image and human annotations is 84.2%, i.e., at least one correctly predicted frame most of the time.
@@
To evaluate the topics output by the model, we use the 19 most frequent topics for our analysis, discarding one of the top 20 topics (Media) because most of its included documents were publication boilerplate rather than article texts. Two of the authors hand-annotate a set of 190 articles (10 per topic), marking whether the model’s topic prediction was acceptable. The overall accuracy for topics predicted by the model is 86% and 87% per annotator. We discard two topics (Business, Environment) with the lowest accuracy scores, leaving us with a final set of 17 topics. We list these, with examples and accuracies, in Appendix Table 7.
@@ 5.1 Generic Framing Analysis
We first analyze the generic frames predictions across the dataset [^4] For framing analysis, we perform additional filtering, excluding articles with “None” or other non-valid frame predictions. We further removed articles less than 100 words and articles on the topics ”sports” and ”media”, as we observed while conducting qualitative analysis that articles on sports only used the Cultural Identity frame in the images and articles with the media topic were of generally low quality with several pages only including text from a paywall, possibly skewing the results. This resulted in a reduced set of  154k articles in  Figure 4. Overall, we can see that there are many more predicted frames for text compared to the images (a mean of 3.6 predicted frames per text and 1.3 predicted frames per image). This is intuitive (and can be seen in the  Figure 1) since the text of the article can more easily express several distinct frames. Looking at their distributions, we can see that in the text of the article, quality of life, crime, and policy appear more frequently, while morality and external regulation & reputation are more rare. In terms of image frame distribution, external regulation & reputation has very low frequency. Contrasting the two modalities, economic and policy frames more frequently appear in text, while capacity & resources and culture are more prevalent in images.
@@
Our results above indicate that frames are used in images and texts in different ways, but how those uses differ is unclear. Why might the same article include a crime frame for its text and a political frame for its image? To explore this question, we perform a lexical analysis of the words used in articles whose image or text used the same frame. We use the “Fightin’ Words” algorithm from Monroe et al. (2008), a comparison metric which takes into account disproportionate numbers of samples as well as rare words, to find the bigrams from the article texts most associated with a single frame for images vs texts. Table 5.1 shows the bigrams sorted by their $z$-scores (prior=$0.01$, frequency5absent5$\geq$ 5≥ 5) for two frames. Qualitatively, we observe that words associated more with image frames tend to be concrete (“ice cream”) and associated with a single meaning more easily recognized to the predicted frame (“police department”). Prior work suggests that more tightly clustered and recognizable images are associated with more concrete topics (Hessel et al., 2018). See the Appendix for a full list of the frames and sorted bigrams.
@@ 5.2 Case Study: Immigration
In Figure 5, we show that articles about immigration use the crime and punishment, external regulation & reputation, and economic framing much more frequently than the images. These articles often focus on deals with other countries, their contribution to the economy, the cost of their deportation, and/or the crimes that they commit. On the other hand, images tend to use the capacity & resources framing, showing migrants in camps, or the public opinion framing, showing people giving speeches. There are also differences across the political spectrum, as can be seen in Figure 8, with the right focusing much more on the security & defense framing compared to sources from the left or center across both modalities. To highlight this further, we show the finer-grained issue specific frames per political leaning in Figure 10. We see clear differences and fine-grained signals about which specific framing publishers leaning to different sides of the political spectrum use. The left and center tend to highlight the humanitarian crisis framing much more, the gap is smaller for right leaning publishers, while on the right, the focus is more on immigrants being an economic burden or a security threat.
Figure 10: Percentage of issue frames used by news publishers in the article text across the political spectrum in their reporting on immigration.

### v3

@@ 1 Introduction
Figure 1: News can be intentionally framed to affect reader perception. Editorial choices decide what is communicated through words and images. Our approach systematically detects this framing.
@@
While visual framing analysis has been explored in communication science and journalism studies (Wessler et al., 2016; Powell et al., 2017), existing studies doing computational framing analysis have largely ignored this crucial aspect (Ali and Hassan, 2022). Further, they have primarily focused on a fixed set of labels for framing analysis, performing prediction in a multi-class setting. This substantially limits the information one can derive from predictions as an article can convey several frames (Figure 1) and fine-grained analysis of framing within a topic necessitates frames specific to a particular issue. Large vision and language models are particularly well suited for conducting this task at scale, considering the semantic understanding embedded into them through large scale pre-training. This led us to our overall research questions: (RQ1) Can LLMs and VLMs reliably detect framing in news articles? (RQ2) Are there differences in framing conveyed through text vs the images? and (RQ3) How do these framings vary across topics and publishers?
Our contributions include the following.
@@
  • We provide a large-scale dataset of 500k U.S.-based news articles for framing analysis, automatically labeled with generic and issue-specific frames, validated through human annotations. [^1] https://huggingface.co/datasets/copenlu/mm-framing
@@ 2 News Framing
Framing Through Visuals
@@
Photographs are an important vehicle of framing as a reader may process textual and visual messages differently. Readers may focus on photographs without also reading an accompanying story Miller (1975) or might select which news story to read depending on the image thumbnail. Images are potent framing tools because they evoke immediate emotional responses, provide contextual cues, and sometimes contradict textual narratives Geise and Xu (2024). For visual data, however, automation is more challenging because computational image analysis often struggles with connotative and symbolic elements that are readily discernible to human annotators Rodriguez and Dimitrova (2011). While some recent advancements in machine learning and computer vision offer promising avenues for automated visual framing analysis, the complexity of symbolic and ideological meanings typically requires human interpretation Coleman (2010).
Automated Framing Analysis
@@
In computational studies, framing has been studied with the help of machine learning methods for content analysis at scale.  Ali and Hassan (2022); Vallejo et al. (2023); Otmakhova et al. (2024) survey these efforts towards on computational framing analysis, focusing on different aspects like methods, varying conceptualisations and inter-disciplinary connections of framings. The most widely used resource is one by  Card et al. (2015), who present the Media Frames Corpus (MFC), a dataset of US based news headlines annotated for generic frames. Most studies use it in a supervised, multi-class prediction setting with MFC frames for analysis of social networks (Mendelsohn et al., 2021) or discussion forums (Hartmann et al., 2019). There are also English-only and multi-lingual SemEval tasks on frame detection (Piskorski et al., 2023a, b; Sajwani et al., 2024). Unsupervised approaches to framing analysis use framing lexica (Field et al., 2018), clustering (Burscher et al., 2016; Ajjour et al., 2019), and topic models (Nguyen, 2015). All of these approaches have focused on texts alone.
Integrative Framing Analysis
Integrative framing analysis is when both images and text are observed separately, but the results are integrated to form a more complete picture of framing analysis (Dan, 2017). Although there is broad agreement that visual and verbal elements should be studied side by side (Coleman, 2010; Rose, 2022), relatively few studies have effectively combined the two. Understanding multimodal framing is important for several reasons: it allows for a more granular examination of media bias, as textual analysis alone may miss the ideological and emotional undertones of visual elements Wessler et al. (2016); Geise et al. (2025). It improves fact-checking and detection of misleading content by identifying inconsistencies between textual claims and their supporting images. It also finds applications in political communication, journalism studies, and public policy through tracing changes in framing over time and across outlets Baumgartner et al. (2008).
@@ 3 Dataset
We crawl news articles along with corresponding images from 28 US-based news agencies, extracting data for a 12 month period between May 2023 and April 2024 using the newsplease library (Hamborg et al., 2017). Our selected news sources reflect the entire political spectrum, based on data from AllSides Media Bias, [^2] https://www.allsides.com/media-bias/ratings an organization that assigns a rating of political leaning to each media outlet. Our list of sources along with corresponding political leanings are listed in Appendix A. We query the publicly available Common Crawl archives [^3] https://commoncrawl.org/news-crawl for the corresponding publishers and extracted each article’s text, headline, publication date, image_urls and other metadata in JSON format. Post scraping, we filter extremely short and long articles, images of logos and other noise. Further details about the filtering process are provided in Appendix C. Our final dataset includes about 500K articles and corresponding images. We show the distribution of the data across the time period of data collection in Figure 3.
@@ 4.1 Model Annotation
We use Large Language Models (LLM) and Vision-Language Models (VLM) to label several aspects of the text and image from news articles, including generic, issue-specific framing and the topic of the article. In Appendix B, we list the different aspects extracted per modality from the news articles. We also generate explanations for the decisions along each prediction for reducing hallucination and facilitating downstream qualitative analysis. For extracting generic frame across both modalities, we use the framing dimensions outlined by Boydstun et al. (2014), as provided in the Media Frames Corpus (Card et al., 2015), which includes 15 generic frames appropriate for analysis of US news. These frames are Economic, Capacity & Resources, Morality. Fairness & Equality, Legality, Constitutionality & Jurisprudence, Policy Prescription & Evaluation, Crime & Punishment, Security & Defense, Health & Safety, Quality of Life, Cultural Identity, Public Opinion, Political, External Regulation & Reputation and Other. The descriptions of each of the frames are provided in Appendix D. Differing from most previous studies on automated framing analysis, which assign a single frame to an article, we conduct the frame extraction in a multi-label setting, i.e., an article or an image can have more than one frame. This setting is much closer to the setup scholars use when conducting qualitative studies on framing (Dan, 2017).
@@ 4.2 Models and Prompts
We use Mistral-7B for text annotations and Pixtral-12B (multi-modal) for image. We use the vLLM library for high-throughput inference, allowing us to conduct analysis on the entire dataset in a few days, demonstrating the scalability of our approach. See Appendix E for more details.
We carefully craft prompts for the extraction of each aspect from the article. For extraction of frames expressed in the articles and images, we experiment with including the Entman and Gamson definitions of framing. We also experiment with short, medium, and long descriptions of each frame and different output formats. We benchmark these different prompts on a validation set and qualitatively examine model decisions and explanations, iteratively improving the prompt by based on categories of errors. For instance, the Pixtral model has a tendency to predict the economic frame every time an entity in professional attire appears in the image, associating professional attire with being wealthy – we instruct the model to avoid such reliance. We report performance on a held-out test set in subsection 4.3. For issue-specific frames, we prompt the model by providing a definition of framing and some examples of issue-specific frames across topics. The task was then to analyse the article and generate an issue-specific frame (described in 2-3 words) w.r.t. the topic of the article. More details are provided in Appendix H. The full list of prompts for the extraction of each aspect is provided in Appendix  1 and  2.
@@ 4.3 Model Evaluation
To assess the quality of output frames from our approach, and compare against prior work, we evaluate our text frame extraction model on the existing large scale benchmark Media Frames Corpus (Card et al., 2015). This corpus provides frame labels by each annotator for over 32,000 articles from US news on topics like immigration, smoking, and same-sex marriage. We take the union of all frame labels assigned by the annotators and assign the top 3 most frequent frames assigned by the annotators as the labels for an article. We run our text framing analysis model on this dataset, providing the article of the text while allowing the model to generate multiple frames per article. We calculate the intersection between the sets of Mistral model annotated frames and human-annotated frames for each article in the dataset. 95.7% of the articles had a non-zero intersection, with at least one overlapping frame label, demonstrating that the model outputs accurate frames in most cases. Our model received a micro averaged F1 score of 0.5, with an averaged precision of 0.42 and averaged recall of 0.62. For prediction across 15 labels, in a multi-label setting, and a task as subjective as framing, we believe the model performs quite well, with several errors being attributed to subjective interpretation. We provide an error analysis of the model with per-label metrics, frequent misclassifications and their examples in Appendix G.
@@
For images, there is no existing benchmark with frame labels. Two of the authors of this study manually annotated 600 images for generic frames across a stratified (time-period and topics) sample of the dataset. We use annotation guidelines released by Mendelsohn et al. (2021), adapting them to the image annotation setting with examples of images for each framing category. We set up a multi-label classification platform (Figure 12) where annotators select one or more frames from the list given an image. To calculate agreement, we compute both Krippendorff’s alpha ($\alpha=0.393$) and mean Jaccard Index (0.614) as measures of inter-annotator agreement. While Krippendorff’s alpha provides label-level reliability, Jaccard Index is particularly well-suited to our multi-label setting, as it evaluates instance-level agreement and gives partial credit when annotators agree on some but not all labels. Framing analysis is a highly subjective task, as is well established by prior work in communication studies as well as NLP (Card et al., 2015). For images, the subjectivity increases substantially, given the limited amount of context available and requiring cultural and conceptual familiarity more so than needed for textual framing analysis (Geise and Baden, 2015). Our agreement scores are thus in line with prior work on framing.
To further minimise the effect of subjectivity of the annotations on our findings, we take the union of the frame annotations by two annotators as our gold set. The model is thus tasked with generating all labels annotated by the different annotators. We calculate the intersection of model predictions and the gold set for each image. The proportion of non-zero intersection instances between the model and human annotations is 84.2%, i.e., at least one correctly predicted frame most of the time, demonstrating the model aligning with the human framing interpretations in a majority of the cases, with the most frequent error being a ‘None’ prediction. We conduct a thorough analysis of frequent mis-classifications by the model in Appendix G.
@@
To evaluate the topics output by the model, we use the 20 most frequent topics for our analysis, discarding one (Media) because most of its included documents were publication boilerplate. Two of the authors hand-annotate a set of 190 articles (10 per topic), marking whether the model’s topic prediction was acceptable. The overlap between the annotators was 83.5%, substantially higher than random chance (50%). When calculated against the labels assigned by each annotator, the overall accuracy for topics predicted by the model was found to be 86% and 87%. We list these, with examples and accuracies, in Table 11.
@@ 5.1 Generic Framing Analysis
We first analyze the generic frames predictions across the dataset [^4] For framing analysis, we perform additional filtering, excluding articles with ‘None’ frame predictions. We further removed articles less than 100 words and articles on the topics ‘sports’ and ‘media’, as we observed while conducting qualitative analysis that articles on sports only used the Cultural Identity frame in the images and several articles with the media topic only had paywall text. This resulted in a reduced set of 154k articles in  Figure 4. Overall, we can see that there are many more predicted frames for text compared to the images (a mean of 3.6 predicted frames per text and 1.3 predicted frames per image). This is intuitive (and can be seen in the  Figure 1) since the text of the article can more easily express several distinct frames. Looking at their relative distributions, we can see that in the text of the article, quality of life, legality, and policy appear more frequently, while morality and capacity & resources are more rare. In terms of image frame distribution, external regulation & reputation has very low frequency, while culture and quality of life are the most frequent. Contrasting the two modalities, economic and policy frames much more frequently appear in text, while capacity & resources and culture are more prevalent in images.
@@
Our results above indicate that frames are used in images and texts in different ways, but how those uses differ is unclear. To explore this question, we perform a lexical analysis of the words used in articles whose image or text used the same frame. We use the “Fightin’ Words” algorithm from Monroe et al. (2008), a comparison metric which takes into account disproportionate numbers of samples as well as rare words, to find the bigrams from the article texts most associated with a single frame for images vs texts. Table 5.1 shows the bigrams sorted by their $z$-scores (prior=$0.01$, frequency5absent5$\geq$ 5≥ 5) for two frames. Qualitatively, we observe that words associated more with image frames tend to be concrete (“ice cream”) and associated with a single meaning more easily recognized to the predicted frame (“police department”). Prior work suggests that more tightly clustered and recognizable images are associated with more concrete topics (Hessel et al., 2018). See the Appendix for a full list of the frames.
@@ 5.2 Case Study: Immigration
In Figure 6, we show that articles about immigration use the crime and punishment, external regulation & reputation, and economic framing much more frequently than the images. These articles often focus on deals with other countries, their contribution to the economy, the cost of their deportation, and/or the crimes that they commit. On the other hand, images tend to use the capacity & resources framing, showing migrants in camps, or the public opinion framing, showing people giving speeches. There are also differences across the political spectrum, as can be seen in Figure 9, with the right focusing much more on the security & defense framing compared to sources from the left or center across both modalities. To highlight this further, we show the finer-grained issue specific frames per political leaning in Figure 10. When looking at frame frequencies normalised by each political leaning, we see clear differences and fine-grained signals about which specific framing publishers leaning to different sides of the political spectrum use. The left and center tend to highlight the humanitarian crisis framing much more, the gap is smaller for right leaning publishers, while on the right, the focus is more on immigrants being an economic burden or a national security threat.
Figure 10: Normalised count of top 10 issue frames used by news publishers in the article text across the political spectrum in their reporting on immigration.